OpenStudy (anonymous):

what is the value of y, if 3y, 4y,and 5y are the exterior angles of a triangle?

OpenStudy (anonymous):

What is always the sum of the exterior angles of a triangle?

OpenStudy (anonymous):

an exterior angle is equal to sum of two interior angles

OpenStudy (anonymous):

Yes, but what is the sum of ALL 3 exterior angles? It always equals a specific number. Maybe draw a few triangles to convince yourself of this? Once you discover this fact, the rest of the problem is easy.

OpenStudy (anonymous):

OK...thank you

OpenStudy (anonymous):

360 =D

OpenStudy (anonymous):

Exterior angle always add to 360

OpenStudy (anonymous):

3y + 4y + 5y = 360 12y = 360 y = 360 / 12 y = 30